# Pay-per-search models are abstention models

### v1

@@ Abstract
LLMs cannot reliably recognize their parametric knowledge boundaries and often hallucinate answers to outside-of-boundary questions. In contrast, humans recognize their limitations and can either seek external help for such questions or abstain. In this paper, we introduce MASH (Modeling Abstention via Selective Help-seeking), a training framework that readily extracts abstentions from LLMs. Our key idea is that any external help-seeking by an LLM, i.e. search tool use, can serve as a proxy for abstention if the external help (search) is appropriately penalized while simultaneously rewarding answer accuracy. MASH operationalizes this idea using reinforcement learning with a pay-per-search reward.
We run experiments on three knowledge-intensive QA datasets. Our results show that MASH  substantially improves upon the selective help-seeking performance of prior efficient search approaches; on multi-hop datasets, MASH improves answer accuracy by $7.6\%$. Furthermore, MASH demonstrates strong off-the-shelf abstention – it can distinguish between unanswerable/answerable questions and selectively generate responses for answerable questions – showcasing behavior analogous to specialized abstention approaches. We emphasize that contrary to prior abstention methods, MASH does not require pre-determining knowledge boundaries to construct training data. Instead, MASH’s abstentions are a by-product of training for the auxiliary selective help-seeking task. Overall, we show that MASH training effectively aligns search tool use with parametric knowledge, which can be successfully leveraged for making abstention decisions. [^1] Code and checkpoints are available at https://github.com/momergul/mash.
@@ 1 Introduction
A reliable AI assistant should recognize its knowledge boundaries – what questions it can and cannot effectively respond to – and act accordingly when a question is outside its boundaries. Conventionally, LLMs learn their knowledge boundaries through alignment by explicitly training for abstention (Yang et al., 2024; Cheng et al., 2024) and calibrated verbalization of uncertainty (Xu et al., 2024b; Stengel-Eskin et al., 2024). These strategies yield improved recognition of capability boundaries but are limited to reducing model errors. The number of questions a model can correctly answer remains unchanged. In this paper, we ask – can we design a training strategy that intrinsically yields an abstention model capable of recognizing its boundaries, while learning techniques that expand its set of answerable questions?
@@
We look at human behavior for inspiration. Humans recognize their limitations and can abstain when asked for knowledge they cannot provide. Alternatively, they can seek outside help which can be used to answer these otherwise unanswerable questions. In this paper, we propose MASH (Modeling Abstention via Selective Help-seeking), a framework that indirectly trains LLMs for abstention by instead training a model to engage in selective help-seeking, i.e. asking for help only when it cannot effectively respond to a query alone.
@@
As a proof of concept, we explore this idea in the context of short-form question-answering tasks. We operationalize help-seeking as invoking a retrieval tool that returns information related to a given query. We train LLMs that selectively seek help (i.e. invoke retrieval) end-to-end with reinforcement learning using a pay-per-search penalty that discounts a correctness reward by the number of searches a model performs. An optimal policy optimizing this reward would, by definition, search only when a question cannot be reliably answered with parametric knowledge. In an inference mode with the same access to search, this model will mirror the above selective search behavior. But more importantly, we can readily elicit abstention decisions from this same model by removing its access to search tools – in that case, any search invocation serves as a proxy for abstention (see Figure 1). MASH, under this framing, effectively trains for two capabilities at the cost of one. Crucially, MASH assumes no privileged information regarding knowledge boundaries like standard abstention approaches (Yang et al., 2024; Cheng et al., 2024; Xu et al., 2024b) or require structured multi-agent interactions (Stengel-Eskin et al., 2024; Eisenstein et al., 2025).
@@
We train MASH models using reinforcement learning with a pay-per-search reward (see Figure 1). However, baseline implementations of this idea (Wang et al., 2025a) result in efficient but sub-optimal search behaviors – models can converge to always searching at least once. To address this, we propose a lightweight synthetic data curation and SFT pipeline that, crucially, assumes no information about the LLM’s parametric knowledge. Instead, it serves to inject diverse, albeit parametrically unaligned, search behavior in LLMs to improve exploration in later RL training. Additionally, we extend the reward formulations of prior work (Wang et al., 2025a) to obtain penalties with harsher levels of severity; this is crucial for extracting good help-seeking behaviors via RL.
@@
We run our experiments on 3 different knowledge-intensive datasets, and evaluate both the selective help-seeking performance with regular inference (w/ access to search) and abstention performance (w/o access to search). Our results show that MASH models substantially outperform previous efficient search baselines (Wang et al., 2025a) at balancing answer accuracy and searches. Notably, on multi-hop datasets, MASH reports a $7.6\%$ accuracy improvement with a better distribution of searches. In fact, this performance is on par with search baselines (Jin et al., 2025) that allow any number of searches (up to a max value) without any penalty. We investigate this further and show that this improvement can be attributed to MASH showcasing a broader range of search strategies, i.e. diversity over number of searches, as a direct result of its training recipe.
@@ 2 MASH: Modeling Abstention via Selective Help-seeking
Figure 1: Overview of MASH’s strategy for eliciting abstractions. Help-seeking LLMs are RL-trained to maximize answer accuracy while minimizing the searches. At inference, this same model is used for abstention by removing search access and treating any search requests as abstention.
@@ 2.1 Abstention Framework
Training Objective  We want the language model $\pi_{\theta}$ to recognize its knowledge boundaries. We posit that we can obtain such a model – without privileged information regarding parametric knowledge boundaries – by training the model to maximize its accuracy while minimizing the number of search requests. Specifically, we optimize the following proxy objective:
$\max_{\theta}~\mathbb{E}_{(q,y)\sim D,\tau\sim\pi_{\theta}(\cdot|q;H)}[r_{acc}(y,\tau)\cdot r_{help}(q,\tau)]-\beta D_{KL}[\pi_{\theta}(\tau|q;H)||\pi_{\theta_{init}}(\tau|q;H)],\vskip-5.69054pt$ (1)
where $D$ is the dataset, $r_{acc}(y,\tau)\in\{0,1\}$ is a binary measure of correctness and $r_{help}(q,\tau)\in[0,1]$ is a multiplicative penalty that assigns a lower value the greater the number of searches in $\tau$. We use reinforcement learning, specially the GRPO algorithm (Guo et al., 2025), to optimize this objective.
Eliciting Abstention from a Selectively Help-Seeking Model  Let $\pi_{\theta^{*}}$ be the optimal policy derived using the above objective. This model will selectively seek help – determine whether to answer a given question $q$ as a function of its expected parametric accuracy and the severity of the $r_{help}$ penalty. We re-frame the goal (and our subsequent evaluations) of this help-seeking model from efficiency, i.e. reducing number of searches, to parametric knowledge alignment, i.e. aligning search behavior with presence or absence of knowledge about a given question in the model’s parameters.
Under this re-framing, we can readily elicit abstentions from a selectively help-seeking model by treating any search invocation as a proxy for abstention. Figure 1 illustrates this abstention framework, which we call MASH: Modeling Abstentions via Selective Help-seeking.
@@ 2.2.1 Initializing $\pi_{\theta}$ w/ Warm-start SFT
RL training to optimize Equation 1 should, in theory, result in a model that selectively seeks help. However, in practice, we find that such training converges to sub-optimal policies – either exhibiting degenerate strategies that always or never search, or failing to learn to use the search tool effectively. In our work, we propose a lightweight and model-agnostic synthetic data generation and finetuning pipeline that results in a substantially better initial policy for subsequent RL training. Our data generation pipeline is designed to encourage diversity in the number of searches in model trajectories. Crucially, it requires no information about model’s parametric knowledge boundaries. In fact, we bake this in explicitly by generating the synthetic fine-tuning dataset using a completely different model with different knowledge boundaries.
Synthetic data generation  Our overall algorithm is outlined in Algorithm 1. For each input question $q$ in the training dataset, we randomly sample a target number of searches $l$ for the associated trajectory and perform constrained decoding with the synthetic data generator $G$ to satisfy this constraint. We sample to generate $l$ consecutive thinking and search steps (appended with retrieved documents from retriever $H(\cdot)$). We achieve this by forcibly appending a <<think>> tag after the initial question and after retrieval outputs, and the <<search>> tag after the end of thinking tag <</think>>. We repeat this $l$ times. We sample multiple such trajectories per question, evaluate each and return a correct trajectory if one exists. Otherwise, we return the trajectory with the shortest answer. Note that this constrained decoding process is only used during synthetic data generation.
A warm start SFT step is also included in recent works’ training pipelines to improve subsequent RL training (Guo et al., 2025; Gandhi et al., 2025; Wang et al., 2025b). However, we highlight one key difference. Contrary to prior works, our warm start process does not not target correctness or alignment with model’s parametric knowledge – the two central goals of MASH. In fact, our synthetic data contains $35\%$ errors with respect to answer correctness and, by design, yields a policy whose search behavior is unaligned with its parametric knowledge (discussed in Appendix A.3). The model learns how and when to use searches during RL training.
Algorithm 1 Warm-Start Trajectory Construction
Datapoint $(q,a^{*})$, generator $G$, retriever $H$, maximum searches $l_{max}$, num samples NN
Datapoint $(q,\tau)$ for SFT
Sample random number of searches $l\sim\{0,\ldots,l_{max}\}$
for $i=1\to N$ do
  Initialize current trajectory $\tau_{i}\leftarrow\emptyset$
  for action in seq do
   Append action start tag $\tau_{i}\leftarrow\tau_{i}~+<\text{action}>$
   Generate action $a\sim G(\cdot|q,\tau_{i})$ until <</action>>
   Append action $a$ to trajectory $\tau_{i}\leftarrow\tau_{i}+a$
   if action == search then
     Retrieve top-$k$ documents $o\sim H(a)$
     Append $o$ to trajectory $\tau_{i}\leftarrow\tau_{i}+o$
return $\tau$
@@ 2.2.2 Reward Formulation
Our reward $r(y,\tau)$ is a product of two terms: $r_{acc}(y,\tau)$, which is a binary correctness reward and $r_{help}(q,\tau)$, which is a search tool penalty. We compute $r_{acc}(y,\tau)$ using exact match.
The form and severity of $r_{help}$ will influence the learned help-seeking behavior. For input question $q$ and $G$ output trajectories $\{\tau_{i}\}^{G}_{i=1}$ sampled during GRPO, let $n$ be the number of search queries in the most efficient and correct trajectory $\tau^{\text{ef}}$ and $m$ be the number of queries in the given trajectory $\tau_{i}$. We want $r_{help}$ to appropriately penalize $\tau_{i}$ if $m>n$. There exists an arbitrarily high number of penalty formulations that satisfy this desiderata; we experiment with three:
@@
$r^{\text{OTC}}_{help}(q,\tau_{i})=\begin{cases}1&\text{if }m=n=0\\ \cos(\frac{m\cdot\pi}{2m+c})&\text{if }n=0\\ \sin(\frac{m\cdot\pi}{m+n})&\text{otherwise}\end{cases},$ (2)
@@
3. OTC-Strict which enforces an extremely strict tool use penalty when $m>n=0$. Note that $n=0$ indicates there is a correct trajectory $\tau^{\text{ef}}$ without any searches. We posit that for these cases, any other trajectory $\tau_{i}$ that uses searches should get a 0 reward under a very strict definition of answerability. Therefore, we set $r^{\text{OTC-St}}_{help}(q,\tau_{i})$ to 0 for such cases. We can use any of the above two reward formulations for when $n>0$, but choose OTC’s sinusoidal function to align with prior work.
@@ 3 Experimental Setup
Datasets and Models   We run our experiments on three knowledge-intensive datasets – the single-hop dataset Natural Questions (NaturalQA) (Kwiatkowski et al., 2019), and multi-hop datasets HotPotQA (Yang et al., 2018) and 2WikiMultiHopQA (2Wiki) (Ho et al., 2020). [^2] We find that the “comparison” and “bridge-comparison” questions comprising in 2WikiMultiHopQA have unbalanced answer distributions (skewed towards “no”). This opens up the possibility of reward hacking by exploiting this dataset property. Therefore, we omit these questions from our training and evaluation. We train and evaluate on each dataset separately; this allows us to evaluate MASH across tasks requiring different search strategies and with different distributions of parametrically answerable questions. We perform all training and evaluation on the Qwen2.5-3B base model (Qwen et al., 2025).We deliberately choose the base model over instruct as the latter has already undergone abstention training although the exact training strategy is unknown; we propose MASH as an alternative. We use the E5 retriever (Wang et al., 2022) and the 2018 Wikipedia dump as our knowledge source (Karpukhin et al., 2020).
@@
Warm-start data generation We follow the strategy outlined in Section 2.2.1 to generate warm-start data for each dataset using Qwen2.5-32B base. This ensures that information about knowledge boundaries is not baked into the SFT training data and that samples follow the prescribed format. For each dataset, we randomly sample $1000$ questions from its training set and set $l_{max}=2$. We select the trajectory for each question from $N=5$ samples. Details can be found in Appendix A.3
@@
3.1 Evaluation Details for Inference Mode I: w/ Search Tools

### v2

# MASH: Modeling Abstention via Selective Help-Seeking

@@ Abstract
LLMs cannot reliably recognize their parametric knowledge boundaries and often hallucinate answers to outside-of-boundary questions. In this paper, we introduce MASH (Modeling Abstention via Selective Help-seeking), a training framework that readily extracts abstentions from LLMs. Our key idea is that any external help-seeking by an LLM, i.e. search tool use, can serve as a proxy for abstention if the external help (search) is appropriately penalized while also rewarding answer accuracy. MASH operationalizes this idea using reinforcement learning with a pay-per-search reward. We run experiments on three knowledge-intensive QA datasets. Our results show that MASH  substantially improves upon the selective help-seeking performance of prior efficient search approaches; on multi-hop datasets, it improves answer accuracy by $7.6\%$. Furthermore, MASH demonstrates strong off-the-shelf abstention performance, showcasing behavior competitive with prior abstention methods that additionally require predetermining model knowledge boundaries to construct training data. Overall, we show MASH training effectively aligns search tool use with parametric knowledge, which can be successfully leveraged for making abstention decisions and efficient search tool use.  [^2] Code and checkpoints are available at https://github.com/momergul/mash.
Machine Learning
@@ 1 Introduction
A reliable AI assistant should recognize its knowledge boundaries, what questions it can and cannot effectively respond to, and act accordingly when a question is outside its boundaries. Conventionally, LLMs learn their knowledge boundaries through alignment by explicitly training for abstention (Yang et al., 2024; Cheng et al., 2024) and calibrated verbalization of uncertainty (Xu et al., 2024b; Stengel-Eskin et al., 2024). These strategies yield improved recognition of capability boundaries but are limited to reducing model errors. The number of questions a model can correctly answer remains unchanged. In this paper, we ask if we can design a training strategy that yields an abstention model capable of recognizing its boundaries, while learning techniques that expand its set of answerable questions?
@@
We look at human behavior for inspiration. Humans recognize their limitations and can abstain when asked for knowledge they lack. Alternatively, they can seek outside help to answer these otherwise unanswerable questions. In this paper, we propose MASH (Modeling Abstention via Selective Help-seeking), a framework that indirectly trains LLMs for abstention by instead training models to engage in selective help-seeking, i.e. asking for help only when it cannot effectively respond to a query alone.
@@
As a proof of concept, we explore this idea in the context of short-form question-answering tasks. We operationalize help-seeking as invoking a retrieval tool that returns information related to a given query. We train LLMs that selectively seek help (i.e. invoke retrieval) end-to-end with reinforcement learning using a penalty that discounts a correctness reward by the number of searches a model performs. An optimal policy optimizing this reward would, by definition, search only when a question cannot be reliably answered with parametric knowledge. In an inference mode with the same access to search, this model will mirror the above selective search behavior. More importantly, we can readily elicit abstention decisions from this same model by removing access to search tools. In that case, any search invocation serves as a proxy for abstention (see Figure 1). MASH, under this framing, effectively trains for two capabilities at the cost of one. Crucially, MASH assumes no privileged information regarding knowledge boundaries like standard abstention approaches (Yang et al., 2024; Cheng et al., 2024; Xu et al., 2024b) or require structured multi-agent interactions (Stengel-Eskin et al., 2024; Eisenstein et al., 2025).
@@
We train MASH models using reinforcement learning with a pay-per-search reward (see Figure 1). However, baseline implementations of this idea (Wang et al., 2025a) result in efficient but sub-optimal search behaviors: models can converge to always searching at least once. To address this, we propose a lightweight synthetic data curation and SFT pipeline that, crucially, assumes no information about knowledge boundaries. Instead, it serves to inject diverse, albeit parametrically unaligned, search behavior to improve exploration in RL training. Additionally, we extend reward formulations of prior work (Wang et al., 2025a) to obtain penalties with harsher levels of severity; we found this crucial for extracting good help-seeking behaviors via RL.
@@
We run our experiments on 3 different knowledge-intensive datasets, and evaluate both the selective help-seeking performance with regular inference (with access to search) and abstention performance (without access to search). Our results show that MASH models substantially outperform previous efficient search baselines (Wang et al., 2025a) at balancing answer accuracy and searches. Notably, on multi-hop datasets, MASH reports a $7.6\%$ accuracy improvement with a better distribution of searches. In fact, this performance is on par with search baselines (Jin et al., 2025) that allow any number of searches (up to a max value) without any penalty. We investigate this further and show that this improvement can be attributed to MASH showcasing a broader range of search strategies, i.e. diversity over number of searches, as a direct result of its training recipe.
@@ 2 MASH: Modeling Abstention via Selective Help-seeking
Figure 1: Overview of MASH’s strategy for abstention. Help-seeking LLMs are RL-trained to maximize accuracy while minimizing searches. At inference, this same model is used for abstention by removing search access and treating any search requests as abstention.
@@ 2.1 Abstention Framework
Training Objective  We want the language model $\pi_{\theta}$ to recognize its knowledge boundaries. We posit that we can obtain such a model, without privileged information regarding parametric knowledge boundaries, by training the model to maximize its accuracy while minimizing the number of search requests. Specifically, we optimize the following proxy objective:
$\max_{\theta}$~ $\mathbb{E}_{(q,y)\sim D,\tau\sim\pi_{\theta}(\cdot|q;H)}[r_{acc}(y,\tau)\cdot r_{help}(q,\tau)]$ (1)
$-\beta D_{KL}[\pi_{\theta}(\tau|q;H)||\pi_{\theta_{init}}(\tau|q;H)],$
where $D$ is the dataset, $r_{acc}(y,\tau)\in\{0,1\}$ is a binary measure of correctness and $r_{help}(q,\tau)\in[0,1]$ is a multiplicative penalty assigning a lower value the greater the number of searches in $\tau$ is. We use reinforcement learning, with the GRPO algorithm (Guo et al., 2025), for optimization.
Eliciting Abstention from a Selectively Help-Seeking Model  Let $\pi_{\theta^{*}}$ be the optimal policy derived using the above objective. This model will selectively seek help as a function of its parametric accuracy for a given question $q$ and the severity of the $r_{help}$ penalty. Particularly, we prove in Appendix A that $\pi_{\theta^{*}}$ will always answer a question $q$ parametrically, without help, if and only if its expected parametric accuracy for $q$ is greater than the maximum attainable expected reward for $q$ when seeking help, which relies on $r_{help}$. This is intuitive, as the optimal policy should choose to seek help for question $q$ only if this resulted in an equivalent or higher reward than answering parametrically.
We re-frame the goal (and our subsequent evaluations) of our help-seeking model from efficiency, i.e. reducing number of searches, to parametric knowledge alignment, i.e. aligning search behavior with presence or absence of knowledge about a given question in the model’s parameters. Under this re-framing, we elicit abstentions from a selectively help-seeking model by treating any search invocation as a proxy for abstention. Following from our theoretical analysis in Appendix A, the optimal policy $\pi_{\theta^{*}}$ would abstain on a question $q$ if and only if its expected parametric accuracy is less than or equal to the reward it would achieve if it were to seek help. Figure 1 illustrates this abstention framework, which we call MASH: Modeling Abstentions via Selective Help-seeking.
@@ 2.2.1 Initializing $\pi_{\theta}$ w/ Warm-start SFT
RL training to optimize Equation 1 should, in theory, result in a model that selectively seeks help. However, in practice, we find that such training converges to sub-optimal policies that either exhibit degenerate strategies that always or never search, or that fail to learn to use the search tool effectively. In our work, we propose a lightweight and model-agnostic synthetic data generation and finetuning pipeline that results in a substantially better initial policy for subsequent RL training. Our data generation pipeline is designed to encourage diversity in the number of searches in model trajectories. Crucially, it requires no information about models’ parametric knowledge boundaries.
Synthetic data generation   Our overall algorithm is a simple constrained decoding procedure. Let $G$ be the synthetic data generation model. For each input question $q$ in the warm-start training dataset, we first randomly sample a target number of searches $l\sim\{0,\ldots,l_{max}\}$ to perform, where $l_{max}$ is a hyperparameter controlling the maximum number of allowed searches. We then manually set the action tag (e.g., <<think>>, <<search>> or <<answer>>) the model should output at the start of each turn given the target number of searches and let the model perform generation normally otherwise. We sample $N$ such trajectories per question, evaluate each and preferentially return a correct trajectory. Note that this constrained decoding process is only used during synthetic data generation. Refer to Algorithm  2 in Appendix B.3 for more details.
A warm start SFT step is also included in recent works’ training pipelines to improve subsequent RL training (Guo et al., 2025; Gandhi et al., 2025; Wang et al., 2025b). However, we highlight one key difference. Contrary to prior works, our warm start process does not target correctness or alignment with models’ parametric knowledge, the two central goals of MASH. In fact, our synthetic data contains $35\%$ errors with respect to answer correctness and, by design through the random sampling of search counts, yields a policy whose search behavior is unaligned with its parametric knowledge (discussed in Appendix B.3). We further mitigate any confounding factors by intentionally using a generator model $G$ different from the model to be trained (as discussed in Section 3). The model learns how and when to use searches during RL training.
@@ 2.2.2 Reward Formulation
Our reward $r(y,\tau)$ is a product of two terms: $r_{acc}(y,\tau)$, which is a binary correctness reward and $r_{help}(q,\tau)$, which is a search tool penalty. We compute $r_{acc}(y,\tau)$ using exact match. The form and severity of $r_{help}$ will influence the learned help-seeking behavior. For input question $q$ and $G$ output trajectories $\{\tau_{i}\}^{G}_{i=1}$ sampled during GRPO, let $n$ be the number of search queries in the most efficient and correct trajectory $\tau^{\text{ef}}$ and $m$ be the number of queries in the given trajectory $\tau_{i}$. We want $r_{help}$ to appropriately penalize $\tau_{i}$ if $m>n$. There exists an arbitrarily high number of penalty formulations that satisfy this desiderata; we experiment with three:
@@
$r^{\text{OTC}}_{help}(q,\tau_{i})=\begin{cases}1&\text{if }m=n=0\\ \cos(\frac{m\cdot\pi}{2m+c})&\text{if }n=0\\ \sin(\frac{m\cdot\pi}{m+n})&\text{otherwise}\vskip-5.69054pt\end{cases},$ (2)
3. OTC-Strict which enforces an extremely strict penalty when $m>n=0$. Note that $n=0$ indicates there is a correct trajectory $\tau^{\text{ef}}$ without searches. For these cases, any other trajectory $\tau_{i}$ that uses searches should get no reward under a very strict definition of answerability. Therefore, we set $r^{\text{OTC-St}}_{help}(q,\tau_{i})$ to 0 for such cases. We can use any of the above two reward formulations for when $n>0$, but choose OTC’s sinusoidal function to align with prior work.
@@ 3 Experimental Setup
Datasets and Models   We run our experiments on three knowledge-intensive datasets: the single-hop dataset Natural Questions (NaturalQA) (Kwiatkowski et al., 2019), and multi-hop datasets HotPotQA (Yang et al., 2018) and 2WikiMultiHopQA (2Wiki) (Ho et al., 2020). [^3] We find that the “comparison” and “bridge-comparison” questions comprising in 2WikiMultiHopQA have unbalanced answer distributions (skewed towards “no”). This opens up the possibility of reward hacking by exploiting this dataset property. Therefore, we omit these questions from our training and evaluation. We train and evaluate on each dataset separately; this allows us to evaluate MASH across tasks requiring different search strategies and with different distributions of parametrically answerable questions. Particularly, NaturalQA and HotPotQA differ in search strategies but have similar distributions of parametrically answerable questions, while 2Wiki is a multi-hop dataset containing a greater proportion of unanswerable questions than both (see Appendix B.4 for further analysis). We perform all training and evaluation on the Qwen2.5-3B base model (Qwen et al., 2025). We deliberately choose the base model over instruct as the latter has already undergone abstention training although the exact training strategy is unknown; we propose MASH as an alternative. We use the E5 retriever (Wang et al., 2022) and the 2018 Wikipedia dump as our knowledge source (Karpukhin et al., 2020).
@@
Warm-start data generation We follow the strategy outlined in Section 2.2.1 to generate warm-start data for each dataset using Qwen2.5-32B base. Choosing a separate model ensures that information about knowledge boundaries is not baked into the SFT training data, while using the base variant avoids abstention behavior present in the instruct model from entering the training data and confounding results. For each dataset, we randomly sample $1000$ questions from its training set and set $l_{max}=2$. We select the trajectory for each question from $N=5$ samples. Details can be found in Appendix B.3
@@
3.1 Details for Inference Mode I: w/ Search Tools